# Simple and Provable Scaling Laws for the Test-Time Compute of Large Language Models

### v2

@@ Abstract
We propose two simple yet principled algorithms that enjoy provable scaling laws for the test-time compute of large language models (LLMs), which require a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation.
  • The first one is a two-stage knockout-style algorithm: given an input problem, it first generates multiple candidate solutions, and then aggregate them for a final output, via a knockout tournament where pairwise comparisons among the candidates are conducted. Assuming that the LLM can generate a correct solution with non-zero probability and do better than a random guess in comparing a pair of correct and incorrect solutions, we prove theoretically that the failure probability of this algorithm decays to zero exponentially or by a power law (depending on the specific way of scaling) as its test-time compute grows.
  • The second one is a two-stage league-style algorithm, where each candidate solution is evaluated by its average win rate against multiple opponents, rather than eliminated upon loss to a single opponent. Under certain technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm, we prove theoretically that the failure probability of the league-style algorithm also decays to zero exponentially as its test-time compute grows.
Through extensive experiments with two challenging benchmarks, namely GPQA and MMLU-Pro, we validate the proposed theories and demonstrate the outstanding scaling properties of both algorithms.
@@ 1 Introduction
Despite the astonishing advancements of large language models (LLMs) in the past few years, they still face challenges with reliability and stability. This hinders their applications in high-stakes scenarios where a problem need to be solved with success probability $99.9\%$ rather than $90\%$. Similarly, in an LLM-based agentic workflow that involves solving many sub-problems, each of them need to be solved with high success probability in order to ensure the success of the complete workflow, since a single error in the process can lead to an incorrect final output. In these and many other similar scenarios, one is willing to boost the success probability by spending more test-time compute on LLM inference. Common approaches include letting the LLM generate a long chain-of-thought trajectory [38, 17, 26, 47], or asking the LLM to iteratively self-verify or self-refine the solutions that it has generated [7, 24, 8, 44, 45]. Another category of approaches is about sampling multiple candidate solutions and then choosing the best one via majority voting [3, 36, 2, 20], ranking with pairwise comparisons [12], or using an external verifier [14, 32] or trained reward model [5, 21, 49].
The primary goal of this work is to provide some theoretical insights into the full potential of scaling up the test-time compute for LLM inference [31, 42, 48, 39, 47, 11]. Throughout this work, we consider a generic problem formulation where an LLM-based algorithm is given an input problem and asked to output a solution. For conceptual simplicity, we evaluate any solution with a binary metric indicating whether it is correct or incorrect [^1] For example, a solution is “correct” if it contains the right reasoning process and final answer to a math problem, or it meets a set of criteria for an open-ended writing task. It is possible to expand the considered scenarios by adopting more general (discrete or continuous) metrics for evaluating a solution, though this is beyond the scope of our current focus.. In this setting, we seek to design practical and principled algorithms that enjoy provable inference scaling laws in the following sense:
@@
We say that an LLM-based algorithm enjoys a provable inference scaling law for a specific input problem if its success probability (with respect to the inherent randomness of the algorithm) in returning a correct solution to the problem can be boosted arbitrarily close to $100\%$ as its test-time compute grows, provided that certain clearly identified assumptions about the problem and the LLM(s) being used are satisfied.222 For clarity, we emphasize that this definition is stated on a per-instance basis and tailored to an individual input problem, e.g., a multiple-choice question from a question-answering benchmark, or a user query saying “please write a Python function for the quicksort algorithm”.
To this end, we first propose a two-stage knockout-style algorithm that enjoys such provable scaling laws, while requiring a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation. Given an input problem, the proposed algorithm first generates multiple candidate solutions, and then aggregates them for the final output, via a knockout tournament where pairwise comparisons among the candidates are conducted. We prove theoretically that its failure probability in solving a specific problem decays to zero exponentially (Theorem 1) or by a power law (Theorem 2) with respect to the total number of LLM calls, depending on the specific way of scaling. These guarantees rely on two natural (and arguably minimal) assumptions: (i) the LLM can generate a correct solution with non-zero probability, and (ii) the LLM can do better than a random guess in choosing the right winner between a pair of correct and incorrect solutions.
We further propose a two-stage league-style algorithm that also enjoys a provable scaling law. Unlike the knockout-style algorithm that eliminates a candidate upon loss to a single opponent, the league-style algorithm evaluates each candidate by its average win rate against multiple opponents. We prove theoretically that its failure probability in solving a specific problem decays to zero exponentially (Theorem 3) as its test-time compute scales up, under technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm.
Our practical implementations of both proposed algorithms are efficient and scalable, with support for parallel and distributed computation. Empirical results with the challenging GPQA [30] and MMLU-Pro [37] benchmarks validate our theories, and demonstrate the outstanding scaling properties of both algorithms.
@@ 2 A two-stage knockout-style algorithm
    Generation. We first generate $N$ candidate solutions, which can run in parallel. In situations where the final answer to the problem contains only a few tokens (e.g., for multiple-choice problems or math calculation), we require that each solution contains a reasoning process or explanation of the rationale that leads to its final answer, which can be elicited by chain-of-thought prompting [38, 17] for example; such information can be useful for enhancing pairwise comparisons in the next stage.
@@
    Aggregation. We aggregate the candidate solutions via a knockout tournament. At each round of the tournament, the candidates are grouped into pairs randomly, and each pair of candidates are compared for $K$ times. The winner of each pair is the one that is favored for more than $K/2$ times; ties are broken randomly. Only the winners will move on to the next round. The final-round winner at the end of this tournament will be the final output of the algorithm.
A visualization of this process can be found in Figure 1. In a minimalistic implementation, both generation and pairwise comparison can be executed with a single black-box LLM (or an ensemble of multiple LLMs).
Figure 1: A visualization of the proposed two-stage knockout-style algorithm (with $N=8$ in this example).
We next introduce some formal notations, followed by our analysis for both success probability and computational efficiency [4, 15] of the proposed algorithm, to be presented in Sections 2.1 and 2.2 respectively. For concreteness and clarity, our analysis is tailored to the aforementioned minimalistic implementation, but it can certainly be extended to more general cases.
@@
We let $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ denote the probability distribution of the output of one LLM call for generating a candidate solution and for comparing a pair of solutions respectively. Given an input problem $x$, the proposed algorithm first samples $N$ independent candidate solutions
$y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$ (1)
during the generation stage. Then, for each pair of candidates $(y,y^{\prime})$ encountered in the knockout stage, the algorithm samples $K$ independent comparison results
$r_{1},\dots,r_{K}\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime}),$ (2)
and identifies the candidate that is favored by the majority of $\{r_{i}\}_{i\in[K]}$ as the winner. Sampling from $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ throughout the algorithm is the sole source of randomness in the following analysis of success probability. The randomness within $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ can originate from LLM decoding with a non-zero temperature, the randomized choice of prompting method or LLM backend for each LLM call, among others.
@@ 2.1 Analysis of success probability
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\text{ is a correct solution})\geq p_{\text{gen}}>0.$ (3)
@@
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp}}>0.5.$ (4)
In other words, we assume that the LLM can generate a correct solution with non-zero probability, and do better than a random guess in comparing a pair of correct and incorrect solutions. Note that $p_{\text{gen}}$ and $p_{\text{comp}}$ are defined with respect to a specific input problem, not for a distribution of problems or a benchmark.
@@
Intuitively, as $N$ (the number of initial candidate solutions) and $K$ (the number of times that each pair of solutions involved in the knockout stage are compared) grow, it becomes more likely that (i) there exist initial candidate solutions that are correct ones, and (ii) they tend to be selected as the winners in pairwise comparisons against incorrect solutions, which together lead to the correctness of the final output of the algorithm. This is formalized in the following theorem.
@@
If Assumption 1 holds for the input problem, then the failure probability of the proposed two-stage knockout-style algorithm with hyperparameters $N$ and $K$ is bounded by
(final output is ${\mathbb{P}}(\text{final output is incorrect})\leq(1-p_{\text{gen}})^{N}+\lceil\log_{2}N\rceil e^{-2K(p_{\text{comp}}-0.5)^{2}}.$
A proof can be found in Appendix A.1. Notably, the failure probability decays to zero exponentially with respect to the hyperparameters $N$ and $K$ that determine the amount of test-time compute. Another way to interpret this theorem is as follows: for a targeted success probability $1-\delta$ (which can be arbitrarily close to 1 as the failure probability $\delta>0$ approaches zero), it suffices to let $(1-p_{\text{gen}})^{N}\leq\exp(-p_{\text{gen}}N)\leq\delta/2$ and $\lceil\log_{2}N\rceil\exp(-2K(p_{\text{comp}}-0.5)^{2})\leq\delta/2$, namely
@@
In other words, the hyperparameters $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $1/p_{\text{gen}}$ and $1/(p_{\text{comp}}-0.5)^{2}$ respectively.
@@
Although Theorem 1 guarantees that an arbitrarily small failure probability can be achieved, it requires the hyperparameter $K$ to be sufficiently large, depending on the value of $p_{\text{comp}}$ that might be unknown a priori in practice. To resolve this concern, we provide an alternative theorem below, which suggests that scaling up $N$ alone is sufficient, even when $K$ is a fixed constant and thus there is still a good chance that the wrong winner is identified when comparing a specific pair of candidates.
To streamline the statement of our theorem, we introduce the notations $\mathcal{M}_{\text{comp},K}$ and $p_{\text{comp},K}$, which generalize $\mathcal{M}_{\text{comp}}$ and $p_{\text{comp}}$ that appear in Eq. (4). Let $\mathcal{M}_{\text{comp},K}$ denote the probability distribution of the comparison result obtained with $K$ independent LLM calls followed by majority voting (with ties broken randomly). Then we have
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp},K}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp},K}>0.5,\quad\text{where}$
$p_{\text{comp},K}=\begin{cases}p_{\text{comp}}\quad&\text{if }K=1,\\ \sum_{\ell=\lceil K/2\rceil}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}\quad&\text{if }K\text{ is odd},\\ \sum_{\ell=K/2+1}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}+\frac{1}{2}{K\choose K/2}p_{\text{comp}}^{K/2}(1-p_{\text{comp}})^{K/2}\quad&\text{if }K\text{ is even}.\end{cases}$
@@
Suppose that Assumption 1 holds and $N$ is a power of 2. Let $p_{i}$ be the probability that a candidate solution at the $i$-th level of the knockout tournament is correct, where $i=0,1,\dots,\log_{2}N$333 In our notations, the zeroth level of the tournament contains the $N$ initial candidates, the first level contains the $N/2$ winners after the first round of pairwise comparisons, and so on. All candidates within the same level of the knockout tournament have the same probability of being a correct one, due to their symmetric roles. Then
$p_{0}=p_{\text{gen}},\quad\text{and}\quad p_{i+1}\geq p_{i}+(2p_{\text{comp},K}-1)(p_{i}-p_{i}^{2})\quad\text{for}\quad i=0,1,\dots,\log_{2}N-1.$
@@
A proof can be found in Appendix A.2. Notably, the linear relationship between $\log_{2}(N)$ and $\log(1/\delta)$ reveals a power-law relationship between the failure probability $\delta$ and the number of candidates $N$.
2.2 Analysis of efficiency
@@
In sum, the total number of LLM calls required by the two-stage algorithm is
$N+K\times\Big{(}\frac{N}{2}+\frac{N}{2^{2}}+\frac{N}{2^{3}}+\dots+2+1\Big{)}\leq N+K\times N=(K+1)\times N,$
whereas the end-to-end latency, if sufficiently many machines are available, is merely
$T_{\text{gen}}+\log_{2}(N)\times T_{\text{comp}},$
where $T_{\text{gen}}$ and $T_{\text{comp}}$ represent the latency of one LLM call for generating a candidate solution and for comparing a pair of solutions, respectively.

### v3

@@ Abstract
We propose two simple, principled and practical algorithms that enjoy provable scaling laws for the test-time compute of large language models (LLMs). The first one is a two-stage knockout-style algorithm: given an input problem, it first generates multiple candidate solutions, and then aggregate them via a knockout tournament for the final output. Assuming that the LLM can generate a correct solution with non-zero probability and do better than a random guess in comparing a pair of correct and incorrect solutions, we prove theoretically that the failure probability of this algorithm decays to zero exponentially or by a power law (depending on the specific way of scaling) as its test-time compute grows. The second one is a two-stage league-style algorithm, where each candidate is evaluated by its average win rate against multiple opponents, rather than eliminated upon loss to a single opponent. Under analogous but more robust assumptions, we prove that its failure probability also decays to zero exponentially with more test-time compute. Both algorithms require a black-box LLM and nothing else (e.g., no verifier or reward model) for a minimalistic implementation, which makes them appealing for practical applications and easy to adapt for different tasks. Through extensive experiments with diverse models and datasets, we validate the proposed theories and demonstrate the outstanding scaling properties of both algorithms.
@@ 1 Introduction
Despite the astonishing advancements of large language models (LLMs) in the past few years, they still face challenges with reliability and stability. This hinders their applications in high-stakes scenarios where a problem need to be solved with success probability $99.9\%$ rather than $90\%$. Similarly, in an LLM-based agentic workflow that involves solving many sub-problems, each of them need to be solved with near-perfect success rate, since a single error in the process can lead to an incorrect final output. In these and many other similar scenarios, one is willing to boost the success probability by spending more test-time compute on LLM inference. One category of methods for this purpose include iterative approaches like generating a sequential chain of thoughts [44, 18, 29, 53, 7, 38] or self-refinement [6, 25, 8, 50, 51]. Another category, which is the focus of this work, is about repeatedly sampling multiple solutions and then aggregating them for the final output [3, 42, 2, 21, 13, 15, 37, 5, 22, 55, 48, 54, 35]. These two categories are complementary and often used together for the best performance [36, 45, 53, 12, 28].
Goal of this work.
We aim to augment the toolkit of LLM inference scaling with practical algorithms and foundational insights. Throughout this work, we consider a generic problem formulation where an LLM-based algorithm is given an input problem and asked to output a solution. For conceptual simplicity, we evaluate any solution with a binary metric indicating whether it is correct or incorrect. We desire algorithms that enjoy provable inference scaling laws in the following sense:
@@
We say that an LLM-based algorithm enjoys a provable inference scaling law for a specific input problem, if its success probability (with respect to the inherent randomness of the algorithm) in returning a correct solution to the problem can be boosted arbitrarily close to $100\%$ as its test-time compute grows, provided that certain clearly identified assumptions about the problem and the LLM(s) being used are satisfied.
Limitations of existing methods.
Strong baseline methods widely adopted in practice may fail (in theory and practice) to achieve this goal, even if a single LLM call already solves the input problem correctly with high or moderate success probability. For example, best-of-N (BoN) sampling with an imperfect verifier might suffer from performance decay as the number of sampled solutions grows because, as explained in Section 5.1 of [5], “the benefits of search are eventually outweighed by the risk of finding adversarial solutions that fool the verifier”. Indeed, prior works highlighted that developing test-time scaling methods that do not rely on perfect verifiers remains an important direction for further research [2, 37]. Majority voting, another strong baseline, might fail for different reasons: even if the LLM has moderate probability, say 45%, of generating a correct final answer, the success probability of majority voting will actually converge to zero as the number of samples grow if there exists an incorrect final answer that a single LLM call generates with probability 46% [3].
Main contributions.
In pursuit of provable inference scaling laws, we propose a two-stage knockout-style algorithm that first generates multiple candidate solutions, and then select one via a knockout tournament where pairwise comparisons among the candidates are conducted. We prove theoretically that its failure probability decays to zero exponentially (Theorem 1) or by a power law (Theorem 2) with respect to the total number of LLM calls, depending on the specific way of scaling. These guarantees rely on two natural (and arguably minimal) assumptions: (1) the LLM can generate a correct solution with non-zero probability, and (2) the LLM can do better than a random guess in choosing the right winner between a pair of correct and incorrect solutions.
We further propose a two-stage league-style algorithm that also enjoys a provable scaling law. Unlike the knockout-style algorithm that eliminates a candidate upon loss to a single opponent, the league-style algorithm evaluates each candidate by its average win rate against multiple opponents. We prove that its failure probability also decays to zero exponentially (Theorem 3), under technical assumptions that are analogous to but more robust than those required by the knockout-style algorithm.
Both proposed algorithms require a black-box LLM and nothing else (e.g., no external verifier or reward model) for a minimalistic implementation, which makes them appealing for practical applications and easy to adapt for different scenarios. Our practical implementations are efficient and scalable, with support for parallel and distributed computation. Empirical results with diverse LLMs (Llama3.1, Qwen2.5, GPT-4o, QwQ-32B) and datasets (GPQA, MMLU-Pro, MATH-500) validate our theories and demonstrate the outstanding scaling properties of both algorithms.
@@ 2 A two-stage knockout-style algorithm
    Generation. We first generate $N$ candidate solutions, which can run in parallel. In situations where the final answer contains only a few tokens (e.g., for multiple-choice problems or math calculation), we require that each solution contains a thinking process, which can be elicited by chain-of-thought (CoT) prompting [44, 18] for example; such information can be useful for enhancing pairwise comparisons in the next stage.
@@
    Aggregation. We aggregate the candidate solutions via a knockout tournament. At each round, the candidates are grouped into pairs randomly, and each pair of candidates are compared for $K$ times. The winner of each pair is the one that is favored for more than $K/2$ times; ties are broken randomly. Only the winners will move on to the next round. The final-round winner at the end of this tournament will be the final output of the algorithm.
Figure 1: A visualization of the proposed two-stage knockout-style algorithm, with $N=8$ in this example.
Figure 1 visualizes this process. For a minimalistic implementation, both stages can be executed with a single black-box LLM (or an ensemble of multiple LLMs). We next introduce some formal notations, followed by our analysis for both success probability and computational efficiency [16, 4] of the proposed algorithm, to be presented in Sections 2.1 and 2.2 respectively.
@@
Let $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ denote the probability distribution of the output of one LLM call for generating a solution and for comparing a pair of solutions respectively. Given an input problem $x$, the proposed algorithm first samples $N$ independent candidate solutions $y_{1},\dots,y_{N}\sim\mathcal{M}_{\text{gen}}(x)$ during the generation stage. Then, for each pair of candidates $(y,y^{\prime})$ encountered in the knockout stage, the algorithm samples $K$ independent comparison results $r_{1},\dots,r_{K}\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})$, and identifies the candidate that is favored by the majority of $\{r_{i}\}_{i\in[K]}$ as the winner. Sampling from $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ throughout the algorithm is the sole source of randomness in the following analysis of success probability. The randomness within $\mathcal{M}_{\text{gen}}$ and $\mathcal{M}_{\text{comp}}$ can originate from LLM decoding with a non-zero temperature, the randomized choice of prompting method or LLM backend for each LLM call, among others.
@@ 2.1 Analysis of success probability
${\mathbb{P}}_{y\sim\mathcal{M}_{\text{gen}}(x)}(y\text{ is a correct solution})\geq p_{\text{gen}}>0.$
@@
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp}}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp}}>0.5.$
In other words, we assume that the LLM can generate a correct solution with non-zero probability, and do better than a random guess in comparing a pair of correct and incorrect solutions. Here, $p_{\text{gen}}$ and $p_{\text{comp}}$ are defined for a specific input problem, not for a distribution of problems or a benchmark.
@@
As $N$ (the number of initial candidate solutions) and $K$ (the number of times that each pair of solutions involved in the knockout stage are compared) grow, it becomes more likely that (1) there exist initial candidate solutions that are correct ones, and (2) they tend to be selected as the winners in pairwise comparisons against incorrect solutions, which together lead to the correctness of the final output of the algorithm. This is formalized in the following theorem.
@@
If Assumption 1 holds for the input problem, then the probability that the proposed knockout-style algorithm returns an incorrect final output decays to zero exponentially with respect to the hyperparameters $N$ and $K$:
${\mathbb{P}}(\text{{failure}})\leq(1-p_{\text{gen}})^{N}+\lceil\log_{2}N\rceil e^{-2K(p_{\text{comp}}-0.5)^{2}}.$
A proof can be found in Appendix A.1. Another way to interpret this theorem is as follows: for a targeted success probability $1-\delta$ (which can be arbitrarily close to 1 as the failure probability $\delta>0$ approaches zero), it suffices to have
@@
In other words, $N$ and $K$ have logarithmic dependence on $1/\delta$, and linear dependence on $1/p_{\text{gen}}$ and $1/(p_{\text{comp}}-0.5)^{2}$ respectively.
@@
Although Theorem 1 guarantees an arbitrarily small failure probability, it requires $K$ to be sufficiently large, depending on the value of $p_{\text{comp}}$ that might be unknown a priori in practice. To resolve this, we provide an alternative theorem suggesting that scaling up $N$ alone is sufficient, even when $K$ is a fixed constant and thus there is still a good chance that the wrong winner is identified when comparing a specific pair of candidates.
To streamline the statement of our theorem, we introduce the notations $\mathcal{M}_{\text{comp},K}$ and $p_{\text{comp},K}$, which generalize $\mathcal{M}_{\text{comp}}$ and $p_{\text{comp}}$ that appear in Assumption 1. Let $\mathcal{M}_{\text{comp},K}$ denote the probability distribution of the comparison result obtained with $K$ independent LLM calls followed by majority voting (with ties broken randomly). Then we have
${\mathbb{P}}_{r\sim\mathcal{M}_{\text{comp},K}(x,y,y^{\prime})}(r\text{ identifies the right winner})\geq p_{\text{comp},K}>0.5,$
where $p_{\text{comp},K}=\sum_{\ell=\lceil K/2\rceil}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}$ if $K$ is odd, and $p_{\text{comp},K}=\sum_{\ell=K/2+1}^{K}{K\choose\ell}p_{\text{comp}}^{\ell}(1-p_{\text{comp}})^{K-\ell}+\frac{1}{2}{K\choose K/2}p_{\text{comp}}^{K/2}(1-p_{\text{comp}})^{K/2}$ if $K$ is even.
@@
Suppose that Assumption 1 holds and $N$ is a power of 2. Let $p_{i}$ be the probability that a candidate solution at the $i$-th level of the knockout tournament is correct, where $i=0,1,\dots,\log_{2}N$111 In our notations, the zeroth level of the tournament contains $N$ initial candidates, the first level contains $N/2$ winners after the first round of pairwise comparisons, and so on. All candidates within the same level of the knockout tournament have the same probability of being a correct one, due to their symmetric roles. Then $p_{0}=p_{\text{gen}}$ and
$p_{i+1}\geq p_{i}+(2p_{\text{comp},K}-1)(p_{i}-p_{i}^{2})\quad\text{for}\quad i=0,1,\dots,\log_{2}N-1.$
@@
A proof can be found in Appendix A.2. The linear relationship between $\log_{2}N$ and $\log(1/\delta)$ reveals a power-law relationship between the failure probability $\delta$ and the number of candidates $N$.
2.2 Analysis of computational efficiency
@@
In sum, the total number of LLM calls required by the two-stage algorithm is $N+K\times\sum_{i}N/2^{i}\leq(K+1)\times N$, whereas the end-to-end latency, if sufficiently many machines are available, is merely $T_{\text{gen}}+\log_{2}(N)\times T_{\text{comp}}$, where $T_{\text{gen}}$ and $T_{\text{comp}}$ represent the latency of one LLM call for generating a candidate solution and for comparing a pair of solutions, respectively.